# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ 3.1.1. Non-oscillatory case
$m\ll\mathcal{X}^{\varepsilon}+q^{2}(CM)^{3}/\mathcal{X}^{1-\varepsilon}\ll q^{2}\sqrt{\mathcal{X}}\ell^{3/2+\varepsilon}$
@@
Observe that the harmonic $e\left({mr^{2}}/{cM\ell q^{2}}\right)$ is automatically flat. Indeed, this is not surprising, since the variables in the Kloosterman sums always lie in the “Linnik range”. It thus suffices to evaluate
@@
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\sum_{r|cM}\frac{1}{r^{2/3}}\,\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\left|\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\right|,$ (3.42)
@@
We will merely consider the $+^{\prime}$ case; the case $-^{\prime}$ can be handled in an entirely analogously manner. Note that the coprimality relation $(r^{2}\overline{cM},\ell q^{2})=1$ has already held. We now resort to the Voronoĭ formula in the ramified case, (2.16), which further converts the double sum $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ essentially into
$\begin{split}&\frac{\mathscr{M}^{1/3}}{\ell q^{2}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{(n/n^{\prime})^{1/3}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\Omega_{\pm}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};{\mathcal{W}_{\pm,r,\ell}}\right)}.\end{split}$
This, however, is controlled by
$\begin{split}&\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)},\end{split}$ (3.44)
@@
Observe that, unlike other scenarios where one only cares about what happens while the parameter $\mathscr{M}$ varying around its height by directly adding a smooth weight after the Cauchy-Schwarz (see, for example, [41, 18]), we need to optimize $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ for any $1\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. Next, we will proceed the argument by distinguishing two different situations.
@@
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{5/7}C^{1+\varepsilon}.$ (3.45)
Here, we have used the estimate $|A_{F}(n,1)|\ll n^{5/7+\varepsilon}$ which follows form the Luo-Rudnick-
Sarnak bound for the generalized Ramanujan conjecture; see, for example, [11, Therem 12.5.1].
When $\mathscr{M}$ is large. Next, let us turn to the case where $\mathscr{M}$ is suitably large, with $\mathscr{M}^{\ast}<\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. We are in a position to rely on the estimates for bilinear Kloosterman sums securing further cancellations. For the sake of the clarity of presentation, one might merely take care of the salient case where $n^{\prime}=1$ (this will be responsible for the dominated contribution); while, for the degenerate case where $n^{\prime}$ is suitably large, our analysis works as well, and the bound that we obtain is dominated by the former (as one will expect). It turns out that the quantity we are really concerned about is the following
$\begin{split}&\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(1,n)}{\sqrt{n}}\,S(\overline{r^{3}}Mc,\pm n;\ell q^{2})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)}.\end{split}$
It is necessary that ${\ell^{3}q^{6}}\gg{(CM)^{3}q^{2}}/\mathcal{X}^{1-\varepsilon}$ for $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, so that the sum over $n$ is not vacuous. Now, an application of Lemma 2.6, with $Z=\mathcal{X}^{\varepsilon}$ and $Z_{1}=Z_{2}=1$, reveals that the display above is dominated in the magnitude by
$\begin{split}&\ll\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\left\{(\ell q^{2})^{2}\sqrt{\frac{Cr}{\mathscr{M}}}+(\ell q^{2})^{3/4}C+\ell q^{2}\right\}\\ &\ll\frac{\sqrt{Cr}(\ell q^{2})^{3/2}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].\end{split}$
@@
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{5/7}C^{1+\varepsilon}+\frac{\sqrt{Cr}(\ell q^{2})^{3/2+\varepsilon}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
@@
$\mathscr{M}^{\ast}=\frac{(\ell q^{2})^{3/2}}{\sqrt{C}r^{3/14}},$
@@
$\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll\ell q^{2}C^{2/3+\varepsilon}r^{4/7}+\frac{\sqrt{CM}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
@@
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\sum_{r|cM}\frac{1}{r^{2/3}}\left(\ell q^{2}C^{2/3}r^{4/7}+\frac{\sqrt{CM}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right]\right).$
This implies that the portion for $r$ being large is less dominated by the focal case of $r=1$. We thus have
@@
For any $\ell\sim L\geq 1$ satisfying that
$L\geq\frac{\sqrt{M}}{q^{2-\varepsilon}},\quad{L^{3}q^{6}}\geq\frac{(CM)^{3}q^{2}}{\mathcal{X}^{1-\varepsilon}},\quad\frac{(Lq^{2})^{3/2}}{\sqrt{C}}\leq\frac{q^{2}(CM)^{3}}{\mathcal{X}^{1+\varepsilon}},$ (3.46)
@@
One notices that the last two conditions in (3.46) requires that
$L\geq\frac{\mathcal{X}^{1/3+\varepsilon}}{q^{8/3}}\quad\text{and}\quad L^{5/6}\geq\frac{\mathcal{X}^{1/6}}{q^{2}M^{1/3-\varepsilon}},\quad L^{3/2}\leq\frac{M^{3}}{q\mathcal{X}^{1+\varepsilon}},$ (3.48)
respectively.
@@ 3.1.2. Oscillatory case
Next, in this part we will focus on the situation where $C$ is suitably small such that $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ in (3.39). Recall (3.40). In this case, the phase of the weight $U_{\ell}(y)$ will be roughly $\sqrt{\mathcal{X}\ell}/CM\gg\mathcal{X}^{\varepsilon}$. Set $A:=4\pi\sqrt{\mathcal{X}\ell}/CM$, and put $\mathscr{J}(y)=$
$U(y)F_{k}^{+}(A\sqrt{y})$ for any $y\in\mathbb{R}^{+}$. To begin with, let us take a glance at the integral
$\int_{\mathbb{R}^{+}}\mathscr{J}(y)e(A\sqrt{y}-3(yx)^{1/3})\mathrm{d}y.$
Notice that $\mathscr{J}$ is compactly supported on $[1/2,5/2]$, and satisfies that $\mathscr{J}^{(j)}\ll_{j,k}1$ for any $j>0$. On finds this integral is negligibly small, unless that $A\asymp x^{1/3}$, and, in which case it is essentially given by
$x^{-1/6}e(-4x/A^{2})\mathscr{J}{\left((2x^{1/3}/A)^{6}\right)}.$
This leads us to the multiple sum
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{mr^{2}}}\,e\left(\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{mr^{2}}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{mr^{2}}{\mathcal{M}}}\right)},$
with $\mathcal{M}\asymp\mathcal{X}^{1/2+\varepsilon}\ell^{3/2}q^{2}$. Based on the analysis in the paragraphs above, it suffices to concentrate on $r=1$ case which, however, is though to be responsible for the dominated contribution. In other words, we are simplified into
@@
Recall that one has the coprimality relation $(cM,\ell q)=1$. Now, if one attempts to apply the reciprocity law and the Voronoĭ, we finally arrive at
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{n^{\prime}|\ell q^{2}}\,\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(cM,\pm n;\ell q^{2}/n^{\prime})\,\widetilde{\mathscr{J}_{\pm,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathcal{M}}{\ell^{3}q^{6}};\frac{c}{C}\right)}$ (3.49)
@@
for some non-zero constant $\rho\in\mathbb{R}$. Observe that the phase is till roughly $\sqrt{\mathcal{X}\ell}/CM$, and the integral $\widetilde{\mathscr{J}_{\pm,\ell}}$ is negligibly small, unless that $\mathcal{M}/(CM\ell q^{2})\asymp x^{1/3}$. This shows that the sum over $n$ is such that
@@
To proceed, one sees that it is necessary that ${\mathcal{X}\ell^{3}q^{4}}/{(CM)^{3}}\gg 1$ for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$; this implies that $\ell\geq\sqrt{M}/q^{2-\varepsilon}$ which is already ensured by (3.26). Returning to (3.49), there are two options for $n^{\prime}$, that is, $n^{\prime}=1$ and $n^{\prime}|\ell,n^{\prime}>1$ ($n^{\prime}=q,q^{2}$ cannot occur since of that $q^{2}>\mathcal{X}^{1+100\varepsilon}\ell^{3}q^{4}/M^{3}$ by the final choice of $\ell$ as presented in (3.54)). We shall merely consider the case of $n^{\prime}=1$. The same method can be adapted to argue another type, the contribution of which is relatively small. As of now, it is clear that
@@
for any $i,j>0$. We now apply Lemma 2.7, with $Z=1,Z_{1}=Z_{2}=\sqrt{\mathcal{X}\ell}/CM$, finding that the expression in (3.49) is dominated by
@@
for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. We thus deduce
@@
For any $\ell\sim L,C\ll\sqrt{\mathcal{X}L}/M^{1+\varepsilon}$ with $L$ satisfying that (3.26), we obtain
@@
Altogether, the various bounds in (3.32), (3.39), (3.47) and (3.50) above will result in
@@
For any $\ell\sim L$ with $L$ satisfying that the constraints (3.25), (3.26) and (3.48), we have
@@
$\phantom{=\;\;}\left.\hskip 61.45805pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{M^{3/2}}+\frac{\sqrt{\mathcal{X}}}{M^{3/2}}+\frac{q}{M}\right).$
@@
In this section, we shall turn to the proof of Corollary 1.3. Indeed, as a direct consequence of Proposition 3.3, by summing over $\ell_{1}$ and $\ell_{2}$ in the amplifier
$|\mathfrak{A}_{f}|^{2}=\sum_{\ell_{1},\ell_{2}\in\mathscr{L}}\alpha_{\ell_{1}}\overline{\alpha_{\ell_{2}}}\lambda_{f}(\ell_{1})\overline{\lambda_{f}(\ell_{2})},$
it turns out that
$L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{M}{L}+{Lq^{5/3}\mathcal{X}^{1/6}}+q\sqrt{{LM}}+{L^{2}\mathcal{X}^{1/4}}\sqrt{q}\right.$
$\phantom{=\;\;}\left.\hskip 114.94893pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{\sqrt{M}}+{\mathcal{X}^{1/4}\ell q^{3/2}}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{\sqrt{M}}+\sqrt{\frac{\mathcal{X}}{M}}+q\right).$ (3.51)

### v4

@@ 3.1.1. Non-oscillatory case
$m\ll\mathcal{X}^{\varepsilon}+q^{2}(CM)^{3}/\mathcal{X}^{1-\varepsilon}\ll q^{2}\sqrt{\mathcal{X}}\ell^{3/2+\varepsilon},$
@@
Observe that the harmonic $e\left({mr^{2}}/{cM\ell q^{2}}\right)$ is automatically flat. Indeed, this is not surprising, since that all of the variables in the Kloosterman sums always lie in the “Linnik range”. It thus suffices to evaluate
@@
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6+\varepsilon}}{CMq^{1/3}}\,\sup_{\begin{subarray}{c}r\leq CM\\ (r,\ell q)=1\end{subarray}}\frac{1}{r^{2/3}}\,\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\left|\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\right|,$ (3.29)
@@
We will merely consider the $+^{\prime}$ case in the following; the case $-^{\prime}$ can be handled in an entirely analogously manner. Observe that the coprimality relation $(r^{2}\overline{cM},\ell q^{2})=1$ has already held by the prior analysis. We now resort to the Voronoĭ formula in the ramified case, (2.10), which further converts the double sum $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ essentially into
$\begin{split}&\frac{\ell q^{2}}{\mathscr{M}^{1/3}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{n^{\prime}n}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\Omega_{\pm}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};{\mathcal{W}_{\pm,r,\ell}}\right)},\end{split}$
where $\overline{r}$ is such that $r\overline{r}\equiv 1\bmod{\ell q^{2}}$. Via Lemma 2.2, this, however, is controlled by the upper-bound of
$\begin{split}&\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell r}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)},\end{split}$ (3.31)
@@
Observe that, unlike other scenarios where one only cares about what happens while the parameter $\mathscr{M}$ varying around its height by directly adding a smooth weight after the Cauchy-Schwarz (see, for example, [43, 20] for relevant details), we need to optimize $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ for any $1\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$ in (3.29). Next, we will proceed the argument by distinguishing two different situations.
@@
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{7/32}C^{1+\varepsilon}.$ (3.32)
Here, we have used the estimate $|A_{F}(m,n)|\ll(mn)^{7/32+\varepsilon}$, which follows from the best record towards to the generalized Ramanujan conjecture for ${\rm{GL}(2)}$ due to Kim and Sarnak; see [29, Proposition 2].
When $\mathscr{M}$ is large. Next, let us turn to the case where $\mathscr{M}$ is suitably large, with
$\mathscr{M}^{\ast}<\mathscr{M}\ll\frac{q^{2}(CM)^{3}}{r\mathcal{X}^{1-\varepsilon}}.$ (3.33)
We are in a position to rely on the estimates for bilinear Kloosterman sums securing further cancellations in (3.31). It turns out that the quantity we are really concerned about is the following
$\begin{split}&\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll\frac{q^{2}(CM)^{3}}{r\mathcal{X}^{1-\varepsilon}}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell r}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)}.\end{split}$
It is necessary that ${\ell^{3}q^{6}}\gg{(CM)^{3}q^{2}}/\mathcal{X}^{1-10\varepsilon}$, with $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, so that the sum over $n$ is not vacuous. Now, an application of Lemma 2.6, with $Z=\mathcal{X}^{\varepsilon}$ and $Z_{1}=Z_{2}=1$, reveals that the display above is dominated in the magnitude by
$\ll\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\sum_{n^{\prime}|\ell q^{2}r}\,\frac{\mathscr{M}^{1/6+\varepsilon}(n^{\prime})^{7/32}}{q\sqrt{\ell r}}\left\{r(\ell q^{2})^{2}\sqrt{\frac{C}{(n^{\prime})^{3}\mathscr{M}}}\right.$
$\phantom{=\;\;}\left.\hskip 150.79968pt+\left(\frac{\ell q^{2}r}{n^{\prime}}\right)^{3/4}(r\overline{r^{2}}M,\ell q^{2}r/n^{\prime})^{1/4}C+\frac{\ell q^{2}r}{n^{\prime}}\right\}$
$\ll\frac{\sqrt{Cr}(\ell q^{2})^{3/2}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
@@
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{7/32}C^{1+\varepsilon}+\frac{\sqrt{Cr}(\ell q^{2})^{3/2+\varepsilon}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
@@
$\mathscr{M}^{\ast}=\frac{(\ell q^{2})^{3/2}}{\sqrt{C}r^{9/32}},$ (3.34)
@@
$\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll\ell q^{2}C^{2/3+\varepsilon}r^{1/32}+\frac{\sqrt{CM}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
@@
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\,\sup_{\begin{subarray}{c}r\leq CM\\ (r,\ell q)=1\end{subarray}}\frac{1}{r^{2/3}}\,\left(\ell q^{2}C^{2/3}r^{1/32}+\frac{\sqrt{CM}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right]\right).$
Observes that the contribution from the portion for $r$ being large (such that $r\gg\mathcal{X}^{\delta},\delta>0$) is much lesser than that from the dominated case of $r=1$ in the expression above. It suffices to concentrate on this case, and we thus are able to deduce
@@
For any $\ell$ given as in (3.8), which satisfies that
$\ell\geq\frac{M^{1/4}}{q^{1-\varepsilon}},\quad{\ell^{3}q^{6}}\geq\frac{(CM)^{3}q^{2}}{\mathcal{X}^{1-\varepsilon}},\quad 1\ll\frac{(\ell q^{2})^{3/2}}{\sqrt{C}}\leq\frac{q^{2}(CM)^{3}}{\mathcal{X}^{1+\varepsilon}},$ (3.35)
@@
Here, the third condition of (3.35) stems from the restriction (3.33) in the dominated case of $r=1$, with the choice of $\mathscr{M}^{\ast}$ as shown in (3.34). One notices that the last two conditions in (3.35) requires that
$L\geq\frac{\mathcal{X}^{1/6+\varepsilon}}{q^{4/3}}\quad\text{and}\quad L\geq\frac{\mathcal{X}^{1/10}}{q^{6/5}M^{1/5-\varepsilon}},$ (3.37)
respectively, upon observing that $L^{2}\leq\ell\leq L^{4}$, whenever $\ell=\ell_{1}\ell_{2}$ or $\ell^{2}_{1}\ell^{2}_{2}$, with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$.
@@ 3.1.2. Oscillatory case
Next, in this part we will focus on the situation where the parameter $C$ is suitably small such that $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ in (3.26). Recall (3.27). In this case, the phase of the weight $U_{\ell}(y)$ will be roughly $\sqrt{\mathcal{X}\ell}/CM\gg\mathcal{X}^{\varepsilon}$. Set $\mathcal{A}:=4\pi\sqrt{\mathcal{X}\ell}/CM$, and put $\mathscr{J}(y)=U(y)F_{k}^{+}(\mathcal{A}\sqrt{y})$ for any $y\in\mathbb{R}^{+}$. To begin with, let us take a glance at the integral
$\int_{\mathbb{R}^{+}}\mathscr{J}(y)e(A\sqrt{y}-3(yx)^{1/3})\mathrm{d}y.$ (3.38)
Notice that $\mathscr{J}$ is compactly supported on $[1/2,5/2]$, and satisfies that $\mathscr{J}^{(j)}\ll_{j,k}\mathcal{X}^{j\varepsilon}$ for any $j>0$. On finds that the integral in $y$ above is negligibly small, unless that $\mathcal{A}\asymp x^{1/3}$, and, in which case it is essentially given by
$x^{-1/6}e(-4x/\mathcal{A}^{2})\mathscr{J}{\left((2x^{1/3}/\mathcal{A})^{6}\right)},$
up to an acceptable error term. It is clear that the integral in (3.38) with the positive sign in the exponential factor is negligibly small, by repeated integration by parts sufficiently many times. Inserting this back into (3.27) gives us the equivalent form of $\Psi(C;\ell,M,q)$ in (3.26) as follows:
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{mr^{2}}}\,e\left(\pm\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{mr^{2}}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{mr^{2}}{\mathcal{M}}}\right)},$ (3.39)
with $\mathcal{M}\asymp\mathcal{X}^{1/2+\varepsilon}\ell^{3/2}q^{2}$ for any $\varepsilon>0$. As hinted by the argument in §3.1.1 above, it suffices to concentrate on the $r=1$ case which, however, is thought to be responsible for the dominated contribution for the display in (3.39). Whilst, for the degenerate case where $r$ is suitably large, our analysis works as well, and the bound that we obtain is dominated by the former (as one will expect). In other words, we are simplified into
@@
Here, as before, it suffices to consider the $+^{\prime}$ case in (3.39), on account of the resemblance of the treatments of both situations, $+^{\prime}$ and $-^{\prime}$. Recall that one has the coprimality relation $(cM,\ell q)=1$. Now, if one attempts to apply the reciprocity law and then the ${\mathrm{GL}}(3)$-Voronoĭ, we finally arrive at
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{n^{\prime}|\ell q^{2}}\,\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(cM,\pm n;\ell q^{2}/n^{\prime})\,\widetilde{\mathscr{J}_{\pm,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathcal{M}}{\ell^{3}q^{6}};\frac{c}{C}\right)},$ (3.40)
@@
for some non-zero constant $\rho\in\mathbb{R}$. Observe that the phase of the integral $\widetilde{\mathscr{J}_{\pm,\ell}}$ is still roughly $\sqrt{\mathcal{X}\ell}/CM$, and it is negligibly small, unless that $\mathcal{M}/(CM\ell q^{2})\asymp x^{1/3}$, by repeated integration by parts sufficiently many times. This, in turn, indicates that the sum over $n$ is such that
@@
To proceed, one sees that it is necessary that ${\mathcal{X}\ell^{3}q^{4}}/{(CM)^{3}}\gg 1$ for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. This implies that $\ell\geq\sqrt{M}/q^{2-\varepsilon}$, which is already ensured by the second pre-condition of (3.2) and the condition we have assumed for the case of $\ell=1$ on Page 17. Recall that $(\ell,q)=1$. Now, returning to (3.40), there are two options for $n^{\prime}$, that is,
  • (1)
    $n^{\prime}=1$;
  • (2)
    $n^{\prime}|\ell,n^{\prime}>1$, with $\ell\geq 2$.
Observe that, here, $n^{\prime}=q\text{ or }q^{2}$ cannot occur, since of that $q^{2}>\mathcal{X}^{1+100\varepsilon}\ell^{3}q^{4}/M^{3}$ by the final choice of $L$ as presented in (3.46) below, upon recalling that $\ell=1,\ell_{1}\ell_{2}$ or $\ell^{2}_{1}\ell^{2}_{2}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, the modulo set defined as in (3.3) before. We will be ready to utilize Lemma 2.7 to the sum in (3.40). It suffices to merely consider Case (1); the contribution from Case (2) would be relatively negligible, upon recalling the remark below the proof of Lemma 2.7. As of now, it is clear that
@@
for any $i,j>0$. We now apply (2.13), with $Z=1,Z_{1}=Z_{2}=\sqrt{\mathcal{X}\ell}/CM$, finding that the expression in (3.40) is dominated by
@@
for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. In other words, we have established the following
@@
For any $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ with $\ell$ being as in (3.8), we obtain
@@
Altogether, the various bounds in (3.12), (3.26), (3.36) and (3.41) above will result in:
@@
For any =$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\},\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, with $L$ satisfying that the constraints (3.1), (3.2) and (3.37), we have
@@
$\phantom{=\;\;}\left.\hskip 59.75095pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{M^{3/2}}+\frac{\sqrt{\mathcal{X}}}{M^{3/2}}+\frac{q}{M}\right).$ (3.42)
@@
In this section, we shall turn to the proof of Corollary 1.3. Recalling the definition of the amplifier $\mathfrak{A}_{f}$ as presented in (3.4), which is dependent on $f_{0}\in\mathcal{B}^{\ast}_{k}(M),$ and the property (3.5), via the positivity of the central value $L(1/2,F\otimes f)$ for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we obtain
$\mathfrak{A}_{f_{0}}L(1/2,F\otimes f_{0})\ll\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\mathfrak{A}_{f}L(1/2,F\otimes f).$
Now, as a direct consequence of Proposition 3.3, by summing over $\ell=1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, it turns out that
$L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{M}{L}+{L^{4}q^{5/3}\mathcal{X}^{1/6}}+qL^{2}\sqrt{{M}}+{L^{8}\mathcal{X}^{1/4}}\sqrt{q}\right.$
$\phantom{=\;\;}\left.\hskip 113.81102pt+\frac{\mathcal{X}^{1/4}L^{7}q^{3}}{\sqrt{M}}+{\mathcal{X}^{1/4}L^{4}q^{3/2}}+\frac{\mathcal{X}^{1/4}L^{5}q^{2}}{\sqrt{M}}+\frac{\sqrt{\mathcal{X}}}{L}+\sqrt{\frac{\mathcal{X}}{M}}+q\right).$ (3.43)